# TalkingHeadBench: A Multi-Modal Benchmark & Analysis of Talking-Head DeepFake Detection

### v2

@@ Appendix F Additional Explainability Results
F.1 Grad-CAM Visualizations for TALL [xu2023tall]
@@
F.2 Grad-CAM Visualizations for DeepFake-Adapter [shao2025deepfake]
@@ Appendix G Demographic Bias Audit
To assess the fairness of the benchmarked detectors and understand biases inherited from the source datasets (FFHQ, CelebV-HQ), we conducted a demographic audit on our test set using classifications from the FairFace model [karkkainen2021fairface]. We analyzed detector performance across three axes: race, gender, and age. Tab 9 summarized below indicate minimal bias along gender lines but reveal notable performance disparities across racial and age subgroups. This analysis highlights an important area for future work in both dataset creation and algorithmic fairness for deepfake detection.

### v3

@@ Appendix F Additional Explainability Results
F.1 Grad-CAM Visualizations for TALL [40]
@@
F.2 Grad-CAM Visualizations for DeepFake-Adapter [32]
@@ Appendix G Demographic Bias Audit
To assess the fairness of the benchmarked detectors and understand biases inherited from the source datasets (FFHQ, CelebV-HQ), we conducted a demographic audit on our test set using classifications from the FairFace model [17]. We analyzed detector performance across three axes: race, gender, and age. Tab 9 summarized below indicate minimal bias along gender lines but reveal notable performance disparities across racial and age subgroups. This analysis highlights an important area for future work in both dataset creation and algorithmic fairness for deepfake detection.